Solve the following equation for the variable of x: (6/x-4)+(7/x+4)=6x/x^2-16
Answer by mananth(16949)

(6/x-4)+(7/x+4)=6x/x^2-16
6(x+4) +7(x-4) = 6x/ x^2-16
(6x+24+7x-28)/ x^2-16 = 6x/ x^2-16
6x+24-7x-28= 6x
6x-4 -7x -6x=0
-7x=-4
x= -4/-7
x= 4/7